### v1

# KernelDNA: Dynamic Kernel Sharing via Decoupled Naive Adapters

Abstract
Dynamic convolution enhances model capacity by adaptively combining multiple kernels, yet faces critical trade-offs: prior works either (1) incur significant parameter overhead by scaling kernel numbers linearly, (2) compromise inference speed through complex kernel interactions, or (3) struggle to jointly optimize dynamic attention and static kernels. We also observe that pre-trained Convolutional Neural Networks (CNNs) exhibit inter-layer redundancy akin to that in Large Language Models (LLMs). Specifically, dense convolutional layers can be efficiently replaced by derived “child” layers generated from a shared “parent” convolutional kernel through an adapter.
To address these limitations and implement the weight-sharing mechanism, we propose a lightweight convolution kernel plug-in, named KernelDNA. It decouples kernel adaptation into input-dependent dynamic routing and pre-trained static modulation, ensuring both parameter efficiency and hardware-friendly inference. Unlike existing dynamic convolutions that expand parameters via multi-kernel ensembles, our method leverages cross-layer weight sharing and adapter-based modulation, enabling dynamic kernel specialization without altering the standard convolution structure. This design preserves the native computational efficiency of standard convolutions while enhancing representation power through input-adaptive kernel adjustments. Experiments on image classification and dense prediction tasks demonstrate that KernelDNA achieves state-of-the-art accuracy-efficiency balance among dynamic convolution variants. Our codes are available at https://github.com/haiduo/KernelDNA 111Preprint. Under review..
1 Introduction
Convolutional Neural Networks [8, 23, 21] have long been built on static convolutional kernels within each layer. While these static kernels enable computationally efficient inference, their fixed, input-agnostic nature inherently limits representation flexibility. Dynamic convolution [28, 4] attempts to address this by adaptively modulating kernels based on input content, but it inevitably scales parameters linearly with the number of kernels, introducing prohibitive memory costs that hinder real-world deployment. Existing model compression techniques, such as pruning [16], quantization [5], and knowledge distillation [6], aim to alleviate parameter growth but require complex multi-stage pipelines: they often depend on fine-tuning pre-trained models with carefully tuned hyperparameters. In contrast, weight-sharing [25, 2] strategies in LLMs demonstrate that parameter redundancy can be exploited for drastic compression without performance loss. Inspired by this, we propose integrating dynamic neural networks’ adaptive capacity with LLM-like weight sharing, unlocking the potential of CNNs through a novel parameter-efficient dynamic modulation framework.
To validate the inherent inter-layer redundancy in conventional CNNs, we visualize the similarity between convolutional kernels across layers using the Linear Centered Kernel Alignment (CKA) index [12]. As shown in Fig. 1, except for a few layers that exhibit lower similarity with others, the majority of layers show high similarity. This observation suggests that compression or transformation can be applied to enable similar layers to share a base layer, thereby reducing the model’s parameters without compromising the architecture. Drawing inspiration from Epigenetics [24]—where phenotypic traits are regulated through non-heritable modifications (e.g., DNA methylation) without altering the underlying genetic code—we design a lightweight adapter module. This module decouples input-dependent dynamic channel attention from kernel-centric spatial modulation, enabling a shared “parent” kernel within a block or stage to generate derived “child” layers via adapter-based transformations. By replacing dense convolutional layers with these adapted child layers, we achieve substantial parameter reduction while preserving performance. Fig. 2 illustrates the distinction between the kernel sharing mechanism in our method and mainstream dynamic convolution approaches.
Previous both paradigms (i.e., Fig. 2 (b) and (c)) share critical limitations: they introduce computationally intensive operations (e.g., dynamic kernel assembly or generation) that degrade inference speed and often cause parameter inflation. In contrast, our kernel-sharing strategy not only reduces parameters but also preserves the efficient inference advantage of traditional convolution by reusing and modulating shared base kernels across layers. Notably, parameter sharing introduces implicit regularization [13], mitigating overfitting risks and even enhancing generalization.
To ensure hardware-friendly efficiency, our adapter integrates three enhanced attention mechanisms: Channel Attention, Filter Attention, and Spatial Attention. These mechanisms allow each convolutional layer to retain distinct characteristics while benefiting from cross-layer (common patterns) captured through shared parameters. Termed KernelDNA, this plug-and-play module functions as an epigenetic “marker”, dynamically tailoring kernel behavior without altering the base parameter structure. It is universally applicable to any standard convolution operator, including regular convolutions, depthwise convolutions, and 11 convolutions.
Extensive experiments on ImageNet-1K and downstream tasks demonstrate that KernelDNA surpasses prior dynamic convolution methods in accuracy while preserving near-original inference speeds (e.g., 7,683 fps vs. ResNet18’s 8,463 fps, 90.7% throughput retention), as shown in Fig. 3. Compared to SOTA variants like ODConv (44.90M params) and CondConv (81.35M params), KernelDNA achieves 1.2–5 parameter reduction (9.24M for ResNet18) with higher accuracy (74.23% vs. 73.97% Top-1). On lightweight models like MobileNetV2, KernelDNA attains 75.51% Top-1 accuracy (vs. KernelWarehouse’s 74.68%) using only 3.14M parameters (vs. KernelWarehouse’s 5.17M), while maintaining 99% of the base model’s throughput (6,260 fps vs. 6,323 fps). These results establish a new Pareto frontier, where KernelDNA uniquely balances parameter efficiency (up to 5 compression), hardware compatibility (near-static latency), and adaptive performance across diverse architectures.
2 Related Work
Dynamic Convolution: Dynamic convolution allows adaptive feature extraction based on input content to address the limited adaptability of static convolution to input variations. Early methods like CondConv [28] proposes conditionally parameterized convolutions, where multiple static kernels are aggregated via input-dependent routing weights. DY-CNNs [4] replaces simple routing with attention mechanisms. DCD [17] leverages kernel matrix decomposition for dynamic channel fusion, eliminating channel-wise dynamic attention. DD [30] decomposed dynamic filters into spatial and channel-specific components, effectively combining position-wise channel and spatial attentions. Recent advances further expanded dynamic convolution’s scope: ODConv [15] introduces dynamic adjustments to kernel spatial size, input/output channel dimensions, and kernel-wise multiplication operations. While improving performance, its reliance on dynamic kernel computations significantly increases FLOPs, degrading inference speed. SD-Conv [10] reduces parameters and FLOPs by sparsifying dynamic kernels, but at the cost of accuracy degradation and unpredictable latency spikes. The SOTA KernelWarehouse [14] assembles convolutional kernels from a fine-grained, static cell-wise kernel bank using input-dependent attention scores. However, the kernel assembly process incurs substantial latency and throughput overhead, making it impractical for resource-constrained devices. In contrast, our KernelDNA leverages a weight-sharing mechanism to preserve the computational efficiency of standard convolution while reducing parameters.
Weight Sharing: The concept of weight sharing originated in NLP, where ALBERT [13] pioneeres cross-layer weight sharing to prevent parameter growth with network depth. This idea later migrated to vision: MiniViT [29] applies weight multiplexing to compress vision transformers while enhancing prediction performance. Recent advances further refine sharing strategies—MPOBERT [20] decomposes parameter matrices into shared local tensors, Dynamic Layer Tying [7] employes reinforcement learning to dynamically select layer-wise sharing relationships during training, and Basis Sharing [25] achieves parameter reuse across layers via singular value decomposition (SVD) for LLM compression. The SOTA Relaxed Recursive Transformers[1] revisits layer tying integrated with Low-Rank Adaptation (LoRA), pushing the frontier of efficient LLM compression. Despite its success in transformers, weight sharing remains underexplored in convolutional networks. To our knowledge, this work is the first to adapt weight-sharing principles to dynamic CNNs, enabling efficient inference and compression.
Visual Attention Module: Visual attention mechanisms enhance the representational power of models by dynamically adjusting feature weights, becoming a core strategy for optimizing CNNs. SE-Net [11] pioneeres channel-wise attention by using global average pooling (Squeeze) and fully connected layers (Excitation) to learn channel weights, amplifying important channels while suppressing redundant ones. CBAM [27] combines channel and spatial attention, generating weights through global pooling and spatial convolutions to calibrate multi-dimensional features. ECA-Net [26] improves SE-Net by replacing fully connected layers with 1D convolutions, avoiding information loss from dimensionality reduction and adaptively adjusting cross-channel interaction coverage. CGC [19] dynamically adjusts convolutional kernel parameters through a context-gating mechanism, integrating input features and global context to generate gating weights, enhancing the contextual relevance of feature representation. In contrast, our KernelDNA adopts a decoupled dynamic-static design, where dynamic Channel Attention operates in parallel with static Spatial/Filter attention, reducing runtime computational dependencies.
3 Preliminaries
Standard Convolution. Given an input feature map with channels and spatial dimensions , the standard convolution operation applies a fixed filter to produce an output feature map . At each spatial location , the convolution is computed as:
where denotes the neighborhood around , and is the bias term. The filter is shared across all spatial locations, making the operation computationally efficient but input-agnostic.
Dynamic Filters. Dynamic filters extend standard convolution by generating input-dependent filters for each spatial location . This allows the convolution to adapt to local features:
However, generating requires significant computational overhead, limiting its application to small networks.
Dynamic Convolution. Dynamic convolution improves efficiency by aggregating a set of static kernels using input-dependent attention weights :
where is computed via a lightweight attention mechanism. This approach enhances representational power with minimal computational overhead.
Although dynamic convolution enhances model capacity by introducing multiple static kernels, it inevitably incurs additional parameters and computational costs, negatively impacting inference efficiency. To address these issues, our proposed KernelDNA tackles the parameter inefficiency of dynamic convolution through a layer-wise weight-sharing mechanism. By integrating lightweight visual adapters, KernelDNA enables shared kernels to retain unique characteristics, effectively replacing dense convolutions in their original positions. The entire process is highly efficient, with minimal impact on throughput.
4 KernelDNA
4.1 Dynamic Kernel Sharing
Prior works attempte to enhance the potential of dynamic convolution by scaling it to finer-grained or larger-scale settings. For instance, the recent SOTA method, KernelWarehouse [14], learns linear combinations of n static fine-grained cell-wise convolutional kernels to explore performance with kernel counts exceeding 100. While it decomposes original kernels into finer-grained components and constructs a shared kernel warehouse, the method dynamically assembles full kernels by weighting cell-level kernels via a contrast-driven attention function based on input features. The reconstructed kernels are then convolved with the input.
Although KernelWarehouse claims to enable cross-layer kernel sharing, it fundamentally follows the traditional kernel bank paradigm—aggregating candidate static kernels via attention-based weighting, albeit at a finer granularity. A detailed implementation is shown in Fig. 4 (a). Despite its improved performance over dense convolutions, KernelWarehouse suffers from a critical flaw: computing attention scores for fine-grained kernels and integrating them into full kernels incur substantial computational overhead. These operations are hardware-unfriendly due to intensive memory access, leading to significantly slower inference speeds compared to standard dense convolutions, thus limiting practical deployment in latency-sensitive scenarios.
In contrast to KernelWarehouse’s fine-grained decomposition and reassembly, our KernelDNA preserves the structure of standard convolutions while introducing layer-wise parameter sharing. Instead of complex computations, KernelDNA employs lightweight adapters to generate layer-specific kernels from shared base parameters, retaining the inference efficiency of standard convolutions. This design enables seamless replacement of traditional convolutions, as depicted in Fig. 4 (b). Where the solid boxes represent static convolutional kernels, which serve as the parent layers in KernelDNA. The dashed boxes denote child layers derived from the parent layers through adapter transformations. These child layers do not introduce additional parameters in storage but are dynamically generated during inference to perform convolution operations.
Furthermore, our adapter integrates both static attention (determined during training) and input-dependent dynamic attention (computed at inference), decoupling purely dynamic attention mechanisms of previous works. This allows flexible adaptation to deployment requirements:
High throughput mode: Disable dynamic attention for minimal latency at a slight accuracy cost.
High accuracy mode: Enable dynamic attention for improved performance with marginal speed trade-offs.
KernelDNA can serve as a plug-and-play replacement for conventional convolution layers, balancing efficiency and adaptability without compromising hardware compatibility.
4.2 Light-weight Modulation Adapter
To efficiently and effectively transform base (parent) kernels into shared (child) kernels, we design a lightweight adapter that incorporates three attention mechanisms to modulate kernel properties across different dimensions: input channel, output channel, and spatial interactions. These mechanisms are termed Channel Attention, Filter Attention, and Spatial Attention, respectively. As illustrated in Fig. 5, Channel Attention is dynamically computed based on input features, while Filter Attention and Spatial Attention are statically precomputed and depend solely on the parent kernel. The static mechanisms can be efficiently implemented in parallel to reduce the computation cost and parameters of kernel banks (KernelWarehouse [14]) and multi-axis attention (ODConv [15]) by 24%–38%.
Notably, the static attention weights can be fused with the parent kernel during initialization to precompute child kernels, trading memory for computational efficiency. Alternatively, child kernels can be dynamically generated at inference time with a minor latency penalty. The mathematical formulation is defined as:
where is the base kernel shared across multiple identical convolutional layers, is the modulated kernel for -th layer, and , , and are three components of -th adapter. The represents the dot product operation.
Channel Attention: In fact, applying Channel Attention to convolutional kernels is equivalent to applying it to feature maps. However, for efficient implementation, we adopt the latter approach. Our implementation resembles SE-Net [11] but incorporates a Batch Normalization (BN) layer after the first fully connected (FC) layer. The default channel reduction ratio is set to 4. The detailed implementation process is as follows:
where the denotes the sigmoid function, denotes represents the global average pooling operation.
Spatial Attention and Filter Attention: Compared to the implementation of Channel Attention, Spatial Attention and Filter Attention are simpler. They are implemented as learnable tensors with shapes and , and are stored as static parameters of the adapter. Once trained, these parameters remain fixed. During inference, they can either be fused during initialization or dynamically computed. The Spatial Attention can modulate the spatial distribution of convolutional kernels, enhancing sensitivity to critical regions (e.g., edges or textures). The Filter Attention can adaptively scale output channels, amplifying discriminative features while suppressing noise. So, our proposed Adapter design enables input-aware kernel modulation while retaining the computational benefits of standard convolutions (Fig. 5(b-d)).
| Purely Dynamic Attention (ODConv) | Decoupled Attention (KernelDNA) |
| - Weight tensor grows with batch size () | - Fixed weight size () |
| - Requires memory-intensive view/reshape operations | - No tensor reshaping needed |
| - parameters per batch | - fixed parameters |
| - Complex grouped convolution | - Standard convolution |
Through analysis, we observe that previous works’ purely dynamic computation of spatial attention and filter attention introduces significant computational overhead. By decoupling these components and computing them in parallel, we achieve both efficiency and expressiveness. This decoupled design inspired the name “DNA” for our adapter. The pseudocode implementation is shown in Algorithm 1.
5 Experiments
In this section, to validate the effectiveness of our proposed KernelDNA, we primarily evaluate image classification accuracy on the ImageNet-1K [22] dataset and downstream task performance (object detection and image segmentation) on the MS-COCO [18] dataset. For simplicity, in the following tables, “KW” denotes the KernelWarehouse method, while “KDNA” represents our proposed KernelDNA. All experiments were conducted on a system equipped with an Nvidia GeForce RTX 4090 GPU and an AMD EPYC 7H12 64-Core Processor CPU. To ensure consistency, latency is reported with an input batch size of 1 and single-threaded CPU execution, while throughput is measured with a batch size of 128. To minimize variance, all results are averaged over 1,000 iterations.
5.1 Image Classification on ImageNet-1K
Experiment Setup: To benchmark against SOTA methods, we evaluate our approach on three representative architectures: standard ResNet18/50 [8], advanced ConvNeXt-Tiny [21], and compact MobileNetV2 [23]. All experiments are implemented using publicly available PyTorch official codebases222We can download related codes from the link., with training strategies aligned with those of KernelWarehouse for fair comparison. In our experiments, b denotes the parameter budget ratio of dynamic convolution methods relative to standard convolution. Notably, ConvNeXt-Tiny employs enhanced training strategies (e.g., advanced data augmentation and learning rate schedules), while other architectures follow conventional training strategies. Detailed configurations include: Optimization is SGD with momentum (0.9) and weight decay (1e-4/4e-5) for ResNet/MobileNetV2, and AdamW with weight decay (5e-2) for ConvNeXt-Tiny. The learning rate is initialized at 0.1 for ResNet, 5e-2 for MobileNetV2, and 4e-3 for ConvNeXt-Tiny, decayed via cosine annealing. Batch size is 256 for ResNet and MobileNetV2, and 1024 for ConvNeXt-Tiny. Training epochs are set to 100/150 for ResNet/MobileNetV2, and 300 for ConvNeXt-Tiny. This setup ensures reproducibility and fairness in comparing parameter efficiency and computational overhead across methods.
Experiment Result: As can be seen from Tab. 1. KernelDNA (KDNA) achieves state-of-the-art accuracy-efficiency trade-offs across diverse architectures (ResNet18/50, MobileNetV2, ConvNeXt-Tiny) while maintaining hardware-friendly inference. Compared to dynamic convolution variants like DY-Conv and ODConv, KDNA reduces parameters by 1.2–5 (e.g., 9.24M vs. 44.90M on ResNet18) with comparable or lower FLOPs (2.28G vs. 1.82G for standard ResNet18), yet outperforms them in accuracy (e.g., 74.23% vs. 73.97% Top-1 on ResNet18). Unlike parameter-heavy methods such as CondConv (81M params), KDNA preserves near-original throughput (7,683 fps vs. ResNet18’s 8,463 fps) and minimizes CPU latency (28.43ms vs. ODConv’s 54.34ms).
| Models | Params | FLOPs | Throughput GPU(fps) | Latency CPU(ms) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | 11.69M | 1.82G | 8463 | 22.16 | 70.25 | 89.38 |
| + SE | 11.78M | 1.82G | 8519 | 23.68 | 70.98 | 90.03 |
| + CBAM | 11.78M | 1.82G | 6023 | 21.25 | 71.01 | 89.85 |
| + ECA | 11.69M | 1.82G | 8091 | 23.54 | 70.60 | 89.68 |
| + CGC | 11.69M | 2.19G | 1537 | 78.92 | 71.60 | 90.35 |
| + DCD | 14.70M | 1.90G | 5305 | 41.43 | 72.33 | 90.65 |
| + CondConv | 81.35M | 1.82G | 280 | 89.90 | 71.99 | 90.27 |
| + DY-Conv | 45.47M | 1.87G | 3308 | 46.15 | 72.76 | 90.79 |
| + ODConv | 44.90M | 1.82G | 1267 | 54.34 | 73.97 | 91.35 |
| + KW | 11.93M | 2.23G | 2272 | 69.62 | 73.67 | 91.17 |
| + KDNA | 9.24M | 2.28G | 7683 | 28.43 | 74.23 | 92.38 |
| ResNet50 | 25.56M | 4.11G | 2432 | 90.77 | 78.44 | 94.24 |
| + DY-Conv | 100.88M | 4.22G | 1161 | 155.63 | 79.00 | 94.27 |
| + ODConv | 90.67M | 4.12G | 560 | 132.39 | 80.62 | 95.16 |
| + KW | 28.05M | 6.15G | 922 | 138.78 | 80.38 | 95.19 |
| + KDNA | 22.78M | 4.33G | 2346 | 93.12 | 80.95 | 95.54 |
| MobiV2() | 1.97M | 0.10G | 12787 | 10.04 | 64.30 | 85.21 |
| + DY-Conv | 4.57M | 0.11G | 5818 | 45.59 | 69.05 | 88.37 |
| + ODConv | 4.44M | 0.11G | 3761 | 62.07 | 70.01 | 89.01 |
| + KW | 2.85M | 0.16G | 3284 | 26.43 | 68.29 | 87.93 |
| + KDNA | 1.88M | 0.11G | 12461 | 25.02 | 69.32 | 89.04 |
| MobiV2() | 3.50M | 0.31G | 6323 | 35.94 | 72.02 | 90.43 |
| + DY-Conv | 12.40M | 0.33G | 3667 | 46.93 | 74.94 | 91.83 |
| + ODConv | 11.52M | 0.32G | 2226 | 59.51 | 75.42 | 92.18 |
| + KW | 5.17M | 0.47G | 2288 | 44.58 | 74.68 | 91.90 |
| + KDNA | 3.14M | 0.31G | 6260 | 35.33 | 75.51 | 92.27 |
| ConvNeXt-Tiny | 28.59M | 4.47G | 1847 | 56.20 | 82.07 | 95.86 |
| + KW | 32.99M | 5.61G | 752 | 224.98 | 82.55 | 96.08 |
| + KDNA | 24.32M | 4.51G | 1739 | 60.12 | 82.76 | 96.34 |
KDNA also surpasses semi-filters sharing approaches like KernelWarehouse (KW) in both efficiency and performance. With 1 parameter budgets, KDNA achieves higher accuracy (75.51% vs. KW’s 74.68% on MobileNetV2) while retaining 98% of the base model’s throughput (6,260 fps vs. 6,323 fps). Its static-dynamic decoupling ensures minimal computational overhead, outperforming KW’s kernel assembly (4.33G FLOPs vs. KW’s 6.15G on ResNet50). These results validate KDNA’s ability to unify parameter efficiency, hardware compatibility, and adaptive feature learning, making it a practical solution for real-world deployment.
5.2 Detection and Segmentation on MS-COCO
Experiment Setup: To further validate the generalization capability of our method, we conduct comparative experiments on downstream tasks, including object detection and instance segmentation, against state-of-the-art dynamic convolution approaches. Specifically, we evaluate on the MS-COCO dataset (118K training and 5K validation images), using Mask R-CNN [9] with ImageNet-1K pre-trained models as the backbone. For fairness, all models adhere to the same training strategies as KernelWarehouse [14], including identical data preprocessing (e.g., random resizing, flipping) and hyperparameter configurations. The evaluation metrics include Average Precision (AP) and its variants (AP50, AP75, APS, APM, APL) for object detection and instance segmentation. All experiments are conducted using the MMDetection [3] framework with the same evaluation settings.
Experiment Result: Tab. 2 compares the performance of our proposed KernelDNA (KDNA) with state-of-the-art dynamic convolution methods on object detection and instance segmentation tasks. Evaluated across diverse backbones (ResNet50, MobileNetV2, ConvNeXt-Tiny), KDNA consistently outperforms existing approaches in multiple metrics, demonstrating superior parameter efficiency.
| Detectors | Backbone Models | Object Detection | Instance Segmentation | ||||||||||
| Mask R-CNN | ResNet50 | 39.6 | 61.6 | 43.3 | 24.4 | 43.7 | 50.0 | 36.4 | 58.7 | 38.6 | 20.4 | 40.4 | 48.4 |
| + DY-Conv | 39.6 | 62.1 | 43.1 | 24.7 | 43.3 | 50.5 | 36.6 | 59.1 | 38.6 | 20.9 | 40.2 | 49.1 | |
| + ODConv | 42.1 | 65.1 | 46.1 | 27.2 | 46.1 | 53.9 | 38.6 | 61.6 | 41.4 | 23.1 | 42.3 | 52.0 | |
| + KW | 41.8 | 64.5 | 45.9 | 26.6 | 45.5 | 53.0 | 38.4 | 61.4 | 41.2 | 22.2 | 42.0 | 51.6 | |
| + KDAN | 42.5 | 66.2 | 46.5 | 28.0 | 46.5 | 54.5 | 39.1 | 61.9 | 41.5 | 23.5 | 42.8 | 52.6 | |
| MobileNetV2 () | 33.8 | 55.2 | 35.8 | 19.7 | 36.5 | 44.4 | 31.7 | 52.4 | 33.3 | 16.4 | 34.4 | 43.7 | |
| + DY-Conv | 37.0 | 58.6 | 40.3 | 21.9 | 40.1 | 47.9 | 34.1 | 55.7 | 36.1 | 18.6 | 37.1 | 46.3 | |
| + ODConv | 37.2 | 59.4 | 39.9 | 22.6 | 40.0 | 48.0 | 34.5 | 56.4 | 36.3 | 19.3 | 37.3 | 46.8 | |
| + KW | 36.4 | 58.3 | 39.2 | 22.0 | 39.6 | 47.0 | 33.7 | 55.1 | 35.7 | 18.9 | 36.7 | 45.6 | |
| + KDAN | 37.3 | 59.8 | 40.5 | 23.0 | 40.7 | 48.5 | 34.8 | 56.5 | 36.8 | 19.5 | 37.8 | 47.2 | |
| ConvNeXt-Tiny | 43.4 | 65.8 | 47.7 | 27.6 | 46.8 | 55.9 | 39.7 | 62.6 | 42.4 | 23.1 | 43.1 | 53.7 | |
| + KW | 44.8 | 67.7 | 48.9 | 29.8 | 48.3 | 57.3 | 40.6 | 64.4 | 43.4 | 24.7 | 44.1 | 54.8 | |
| + KDAN | 45.0 | 67.9 | 49.2 | 30.2 | 48.7 | 58.0 | 41.0 | 65.1 | 43.6 | 24.8 | 44.5 | 55.2 | |
Notably, KDNA surpasses 4-parameter methods like ODConv (detection : 42.5 vs. 42.1 on ResNet50) and outperforms KernelWarehouse (KW) under equal parameter constraints (detection : 37.3 vs. 36.4 on MobileNetV2). Its decoupled dynamic-static design enhances multi-scale adaptability, improving small-object detection () by +0.8 and large-object segmentation () by +0.4 over KW on ConvNeXt-Tiny. Crucially, KDNA maintains hardware efficiency via static attention pre-fusion, exceeding the throughput of 4-parameter methods while enabling plug-and-play deployment. This demonstrates its superior balance of accuracy, parameter efficiency, and practicality for real-world vision tasks.
5.3 Ablation Study
To comprehensively validate the advantages of our proposed dynamic kernel-sharing mechanism over standard convolution and the effectiveness of the three visual attention modules, we conduct a series of ablation experiments on the ImageNet-1K dataset.
Settings of “parent” and “child” convolutional layers. Although previous experiments have demonstrated the effectiveness of our proposed dynamic kernel-sharing mechanism, the optimal number and placement of shared kernels (“child”) and base kernels (“parent”) require further exploration. Let the shared kernels be denoted as “S” and the base kernels (full kernels) as “F”. Subscripts of “S” and “F” indicate parent-child relationships. Considering that convolutional layers within the same block (typically with consistent kernel sizes) are more amenable to parameter sharing, our experiments focus on sharing and extending kernels across all blocks within the same stage. The results are shown in Tab. 3.
| Models | Parent-Child | Params | Throughput GPU(fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | FF-FF(original) | 11.69M | 8463 | 70.25 | 89.38 |
| FF-FFF | 14.64M | 7944 | 73.41 | 91.37 | |
| FF-FF-FF | 17.59M | 7496 | 74.07 | 92.93 | |
| FS-FSF | 9.07M | 7814 | 73.95 | 92.61 | |
| FS-SSF | 6.29M | 7756 | 73.42 | 91.64 | |
| FS-SF-SF | 9.24M | 7683 | 74.23 | 92.38 |
From the results, it is evident that the placement of parent and child kernels is crucial for model performance. Using preceding layers as child kernels for a parent layer yields better results than using subsequent layers. Additionally, as the number of shared child kernels increases, the model’s performance gains diminish and may even degrade. Thus, simply scaling up the number of child kernels has an upper limit. However, scaling up the number of parent kernels gradually improves performance, albeit at the cost of increased parameters. Based on these experiments, we find that adopting the “FS-SF-SF” sequence and quantity within the same block achieves the best trade-off between inference efficiency and performance. The “-” indicates the position for the residual summation.
Comparison of methods to increase convolutional layers. To validate the advantages of shared child kernels generated from parent kernels via lightweight adapters over simply adding more convolutional layers or directly copying parent kernels, we conduct experiments with three approaches:
Expanding Parent Kernels: Adding more static convolutional layers.
Direct Copying: Creating shared child kernels by directly duplicating parent kernels.
Adapter-Based Generation: Producing shared child kernels through lightweight adapters (our proposed method).
| Models | Increase layers | Params | Throughput GPU(fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | Original | 11.69M | 8463 | 70.25 | 89.38 |
| Expanding Parent | 14.64M | 7944 | 73.41 | 91.37 | |
| Direct Copying | 8.74M | 7460 | 69.85 | 89.41 | |
| Adapter-Based | 9.24M | 7683 | 74.23 | 92.38 | |
| ResNet50 | Original | 25.56M | 2432 | 78.44 | 94.24 |
| Expanding Parent | 26.67M | 2361 | 80.23 | 94.83 | |
| Direct Copying | 22.55M | 2355 | 75.21 | 92.87 | |
| Adapter-Based | 22.78M | 2346 | 80.95 | 95.54 |
The results are presented in Tab. 4. From the experimental outcomes, it is evident that our adapter-based approach achieves a superior trade-off between performance and efficiency compared to the other two methods.
Three attention components analysis. Our comparative study on different methods of expanding convolutional layers demonstrates that sharing parameters via a “bridge” (i.e., the adapter) is critical, yet the design of this bridge is equally pivotal. In this work, we decouple visual attention mechanisms across dimensions into dynamic and static components, addressing channel-wise, spatial-wise, and filter-wise kernel modulation. Compared to prior works like ODConv, which employs additional dynamic and sequential kernel attention, our implementation is simpler and more efficient.
| Models | Channel | Spatial | Filter | Params | Throughput GPU(fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | ✗ | ✗ | ✗ | 8.74M | 7460 | 69.85 | 89.41 |
| ✓ | ✗ | ✗ | 9.24M | 7695 | 71.82 | 90.72 | |
| ✗ | ✓ | ✗ | 8.74M | 7732 | 70.90 | 90.06 | |
| ✗ | ✗ | ✓ | 8.75M | 7786 | 70.29 | 89.29 | |
| ✓ | ✓ | ✗ | 9.24M | 7485 | 72.68 | 91.54 | |
| ✓ | ✗ | ✓ | 9.24M | 7499 | 71.98 | 91.05 | |
| ✗ | ✓ | ✓ | 8.75M | 7606 | 71.01 | 90.23 | |
| ✓ | ✓ | ✓ | 9.24M | 7683 | 74.23 | 92.38 |
To further validate the effectiveness of the three attention components, we conduct ablation studies (see Tab. 5). The results show that all three attention mechanisms contribute positively to performance. Notably, Channel Attention delivers the most significant accuracy gains, albeit with a slight impact on inference efficiency. In contrast, Spatial Attention and Filter Attention achieve moderate performance improvements without introducing additional computational overhead. These findings collectively confirm the efficacy of our lightweight adapter’s tripartite attention design.
6 Discussion
6.1 Why Kernel Sharing is Effective?
While experiments validate the effectiveness of KernelDNA across multiple vision tasks, further exploration into its underlying mechanisms is critical. We conduct Linear CKA analysis on ResNet18 trained with KernelDNA under optimal configurations on ImageNet-1K (Fig. 6). The left subplot shows the similarity grid of the original ResNet18 layers, while the middle subplot represents KernelDNA’s results, where parameter-sharing replaces or extends original convolutional layers, resulting in slightly more layers. The right subplot compares layer-wise similarity between KernelDNA and the original model. We observe that:
Preservation of Structural Patterns: KernelDNA maintains the original model’s global similarity distribution—the last two layers and a few intermediate layers exhibit lower similarity (indicating task-specific representation learning), while other layers remain highly similar.
Block-Wise Specificity: KernelDNA retains intra-block commonalities (shared patterns) between parent and child layers while introducing inter-block heterogeneity (distinct adaptations), aligning with principles of epigenetic regulation in feature learning.
Parameter Sharing Validity: High similarity between specific layers of KernelDNA and the original model confirms that shared parameters preserve critical learned features, demonstrating the efficacy of our mechanism.
This analysis bridges biological inspiration with neural architecture design, showing how KernelDNA balances redundancy reduction with task-specific adaptability, offering a principled framework for efficient dynamic networks.
6.2 Do Adapters Really Learn the Differences?
To further validate whether our proposed “DNA” adapter—equipped with attention mechanisms—genuinely learns diversity across shared child layers, we visualize the spatial attention features, as shown in Fig. 7. The figure compares three child layers derived from the third parent layer and three from the fourth parent layer. The results demonstrate that child layers from the same parent exhibit distinct variations, confirming that our adapter preserves shared commonality features from parent kernels while introducing differentiated diversity among child layers. This aligns with our hypothesis that the adapter balances inheritance and specialization, ensuring both parameter efficiency and expressive power.
7 Limitations
While our proposed kernel-sharing strategy achieves significant parameter reduction, it does not reduce computational costs (i.e., FLOPs) and may introduce marginal additional computational. Though this trade-off might make it less ideal for compute-constrained devices, our method remains broadly applicable given that modern hardware platforms are predominantly memory-bound rather than computation-limited. By prioritizing parameter efficiency without compromising the inference speed of standard convolutions, KernelDNA aligns with the practical demands of most real-world deployment scenarios.
8 Conclusion
KernelDNA introduces a lightweight convolution plugin that addresses the parameter and efficiency limitations of dynamic convolutions. By leveraging cross-layer weight sharing and adapter-based modulation, it combines dynamic kernel adaptation with standard convolution’s efficiency, reducing parameters while enhancing performance. Experiments show state-of-the-art accuracy-efficiency trade-offs on image classification and dense prediction tasks, outperforming existing methods. KernelDNA maintains hardware-friendly inference speeds (e.g., 90–99% of base models’ throughput) with 1.2–5 fewer parameters than SOTA variants, offering a practical solution for adaptive CNNs without accuracy compromises.